OpenStudy (anonymous):

The Johnsons need a new living room couch. The couch they like costs $2,700. The salesperson told the Johnsons that the couch will be on sale next month for 20% off. If the sales tax rate in their city is 7.5%, how much will the Johnsons save including sales tax, if they wait until the couch goes on sale?

OpenStudy (anonymous):

@jim_thompson5910 help pls!

OpenStudy (anonymous):

Savings = Price if they pay now - Price if they wait = 1.075 * 2700 - 1.075 * (0.8 * 2700)

OpenStudy (anonymous):

You can add on a 7.5% sales tax by just multiplying by 1.075 To reduce something by 20%, just multiply it by (1 - 0.2) or 0.8

OpenStudy (anonymous):

wait what im confused @Mertsj

OpenStudy (mertsj):

Would you agree that right now they have to pay 100% of 2700 plus the sales tax?

OpenStudy (anonymous):

yes @Mertsj

OpenStudy (anonymous):

= 1.075 * 2700 - 1.075 * (0.8 * 2700) = $580.50

OpenStudy (mertsj):

So since the sales tax is 7.5%, right now they have to pay 107.5 percent of 2700 or 1.075(2700) which is $2902.50

OpenStudy (anonymous):

is it D

OpenStudy (mertsj):

Next month the couch is 20% off which means they DON'T have to pay 20%, which means they DO have to pay 80%

OpenStudy (anonymous):

Yes. And the amount they pay if they wait is 1.075 times (80% of 2700) which is 0.86 * 2700 = 2322

OpenStudy (mertsj):

No it is not D . The question asks how much will they SAVE?

OpenStudy (anonymous):

It's C...just as I calculated above

jimthompson5910 (jim_thompson5910):

If you do not wait for the sale, then you simply pay the $2,700 plus 7.5% of that (add on the sales tax) 2700 + 0.075*2700 = 2,902.50 So if you don't wait for the sale, then you pay a total of $2,902.50 ---------------------------------------------------------------------------- Now let's say you wait for the sale of 20% off 20% of $2,700 = (20/100)*2700 20% of $2,700 = 0.20*2700 20% of $2,700 = 540 Taking 20% off of $2,700 means you subtract off $540 2700 - 540 = 2160 Now add on 7.5% sales tax 2160 + 0.075*2160 = 2322 So you will pay $2,322 if you wait next month for the 20% off sale. ---------------------------------------------------------------------------- The difference in those final prices is: 2902.50 - 2322 = $580.50

OpenStudy (anonymous):

Thx i got it! @jim_thompson5910

OpenStudy (anonymous):

And all of that can be simplified to Savings = Price if they pay now - Price if they wait = 1.075 * 2700 - 1.075 * (0.8 * 2700) = 580.5